OpenStudy (rock_mit182):

Solve the following proportion for . Round your answer to the nearest tenth.

OpenStudy (rock_mit182):

\[\frac{ v }{ 10 }=\frac{ 17 }{ 3}\]

OpenStudy (rock_mit182):

v = 170/3

OpenStudy (rock_mit182):

56.6 --> 56.7

OpenStudy (rock_mit182):

the final answer is right ?

OpenStudy (mathlegend):

Yes

OpenStudy (rock_mit182):

thanks brother !